OpenStudy (hannahduhe):

The triangles are congruent. Find the length of each hypotenuse. A.3 B.5 C.17 D.22

OpenStudy (kinged):

Equate the two equations, solve for x, and substitute x in any hypotenuse equation.

OpenStudy (hannahduhe):

is it B

OpenStudy (kinged):

Did you first figure out y by equating its equations?

OpenStudy (hannahduhe):

no

OpenStudy (hannahduhe):

how will the equation look

OpenStudy (kinged):

First, equate\[y+3\]and\[2y+1\]to solve for y.

OpenStudy (hannahduhe):

i got a decimal

jhonyy9 (jhonyy9):

y +3 = 2y +1 y = ? this is an integer

OpenStudy (hannahduhe):

is it 3

OpenStudy (phi):

y +3 = 2y +1 add -1 to both sides (that means write -1 on both sides) can you do that ?

OpenStudy (hannahduhe):

yea so it would look like -1y+3=2y+1-1

OpenStudy (phi):

ok on the right side. on the left side we are "adding" the -1 so it looks like this y +3-1 = 2y +1 -1

OpenStudy (phi):

on the right side we can simplify 1-1 to 0 on the left side we simplify 3-1

OpenStudy (amypurplediaz28):

i dont mean to be rude but if someone may please help me with my question i only have 15 more miutes to answer or i will fail.

OpenStudy (phi):

you should get y +2 = 2y now add -y to both sides

OpenStudy (hannahduhe):

so y+2-y=2y-y right

OpenStudy (phi):

yes. on the left side you have 1 y and take away 1 y. that means 0 y's on the left side on the right side you have 2 y's take away 1 y. how many y's on the right side ?

OpenStudy (hannahduhe):

3 right

OpenStudy (phi):

2 y's take away 1 y is ?

OpenStudy (hannahduhe):

1

OpenStudy (phi):

not just 1. 1y or y

OpenStudy (phi):

so y+2-y=2y-y becomes 2 = y which means y is 2

OpenStudy (phi):

now we use that info to find the hypotenuse. they tell us both triangles are the same size and shape, so the hypotenuses are equal in other words 5x+7 = 6x+2y we can "replace" the y with 2 5x+7 = 6x+2*2 or 5x+7 = 6x+4 I would now add -5x to both sides

OpenStudy (hannahduhe):

ita A 3

OpenStudy (phi):

they want the length of the hypotenuse (longest side). we know that the short side is y+3 which is 5 (because y is 2) so we know the hypotenuse has to be bigger than 5 (and so not 3)

OpenStudy (hannahduhe):

oh k so x is 3 thats what i figured out just now

OpenStudy (phi):

so use x=3 to find the length of the side

OpenStudy (hannahduhe):

ahhh its 22

OpenStudy (phi):

yes

OpenStudy (hannahduhe):

thanks for all your help

OpenStudy (phi):

yw